A commercial bank has no excess reserves until a depositor places $2,000 in cash in the bank. The reserve ratio is 10 percent. The bank then lends $1,500 to a borrower. As a consequence of these transactions, the bank's excess reserves are

A.not affected.
B.increased by $200
C.increased by $300.
D.increased by $500.
Answer #1

Answer : The answer is option C.

Based on given information, the bank's excess reserves occurs when $2,000 is deposited in the bank as a form of cash. The reserve ratio = 10% = 0.1

Bank's reserve = Deposit amount * Reserve ratio = 2000 * 0.1 = $200.

Bank lends $1500.

So, bank's excess reserve = Deposit amount - Bank's reserve - Bank's lending amount = 2000 - 200 - 1500 = $300

Therefore, option C is correct.